OpenStudy (anonymous):

I'm willing to fan and medal. A pair of equations is shown below. 3x − y = 9 y = −2x + 11 If the two equations are graphed, at what point do the lines representing the two equations intersect? (4, 3) (3, 4) (9, 11) (11, 9) Just a bit confused ;-;

OpenStudy (lynfran):

3x-(-2x+11)=9 3x+2x-11=9 5x=9+11 5x=20 x=20/5 x=??

OpenStudy (lynfran):

whats 20/5=??

OpenStudy (anonymous):

Four.

OpenStudy (lynfran):

good x=4 so when x=4 y=-2(4)+11 y=-8+11 y=??

OpenStudy (anonymous):

19 ofc

OpenStudy (lynfran):

no -8+11=?? it can also be written as 11-8...??

OpenStudy (anonymous):

Whoops, didn't see the negative sign. That would be three. THey would intersect at (4, 3) correct?

OpenStudy (lynfran):

correct